OpenStudy (anonymous):

Solve the following system of equations. x + 3y + 2z = –16 2x – y + 2z = –15 2x – 2y – z = 1 (–3, –1, –5) (–3, 5, –2) (2, –4, –3) (4, 4, –1)

OpenStudy (ivettef365):

you can find answer on this website http://calculator.tutorvista.com/system-of-equations-solver.html

OpenStudy (texaschic101):

x + 3y + 2z = - 16 -->(-2)x + 3y + 2z = - 16 2x - y + 2z = - 15 ---------------- -2x - 6y - 4z = 32 (result of multiplying by -2) 2x - y + 2z = - 15 ----------------add -7y - 2z = 17 2x - y + 2z = - 15 -->(-1)2x - y + 2z = - 15 2x - 2y - z = 1 ---------------- -2x + y - 2z = 15 (result of multiplying by -1) 2x - 2y - z = 1 ----------------add -y - 3z = 16 -7y - 2z = 17 -y -3z = 16 -->(-7)-y - 3z = 16 -------------- -7y - 2z = 17 7y + 21z = - 112 (result of multiplying by -7) ---------------add 19z = - 95 z = - 5 -y - 3z = 16 -y - 3(-5) = 16 -y + 15 = 16 -y = 16 - 15 -y = 1 y = -1 x + 3y + 2z = - 16 x + 3(-1) + 2(-5) = - 16 x - 3 - 10 = - 16 x - 13 = - 16 x = - 16 + 13 x = - 3 check... 2x - 2y - z = 1 2(-3) - 2(-1) - (-5) = 1 -6 + 2 + 5 = 1 -6 + 7 = 1 1 = 1 (correct) ANSWER : (-3,-1,-5)